# Training LLMs for Generating IEC 61131-3 Structured Text with Online Feedback

### v1

@@ IV-A Performance Metrics
To quantify the performance of the pipeline, we define evaluation metrics appropriate for the experts $\mathcal{E}_{\kappa}$ and $\mathcal{E}_{\psi}$.
@@
    Compilation, $P(\mathcal{E}_{\kappa}(C_{i})=1)$ – The empirical probability that a code segment $C_{i}$ successfully compiles. This metric reflects the likelihood that the code passes syntax and type-checking stages without errors.
@@
    Semantic correctness, $P(\mathcal{E}_{\psi}(C_{i})=1)$ – The empirical probability that $C_{i}$ is semantically accurate. The primary criterion is the quality of the implementation and its alignment with the intended purpose $I_{i}$.
@@
    Joint semantic correctness and compilation rates, $P(\mathcal{E}_{\psi}(C_{i})=1\land\mathcal{E}_{\kappa}(C_{i})=1)$ – The empirical joint probability that $C_{i}$ is both syntactically valid and semantically correct. This metric reflects the overall success in passing both compilation and semantic validation processes.
These customized metrics provide evaluation for both the quantitative and qualitative aspects of the pipeline.
SYSTEM: “You are an expert in PLC Structured Text (ST) programming as per IEC 61131-3 standards. Your task is to evaluate the given code based on its logical, semantic, and syntactic correctness. You can tolerate minor mistakes or errors. IMPORTANT: Use the following format for your responses: “[0] [1] [1]”, where 0 indicates a negative evaluation and 1 indicates a positive evaluation. IMPORTANT: Ensure that the number of brackets matches the number of given implementations. IMPORTANT: Provide the evaluations strictly in the specified format without any additional text, code, or formatting. INSTRUCTION: Given implementations are separated by ’====================’.” USER: “Great! Here is another problem and its corresponding implementations to evaluate. Problem: $I_{i}$ Implementations: $\mathcal{\hat{C}}_{i}=\{\hat{C}_{1},\hat{C}_{2},\ldots,\hat{C}_{k}\}$
Figure 2: The expert system prompt combines the user prompt, containing the specific intent $I_{i}$, along with its associated synthetic code subset $\mathcal{\hat{C}}_{i}=\{\hat{C}_{1},\hat{C}_{2},\ldots,\hat{C}_{k}\}$.
@@ IV-B Experimental Setup
The baseline model is based on the Phi-3 LLM architecture [4], specifically the version with 14 billion parameters (Phi-3 14B). To maintain its optimal performance, no quantization techniques were applied. The model was reparameterized using SFT with a synthetic dataset generated in advance. This dataset was derived from the APPS dataset [14], which contains prompt-code pairs. The implementations were transpiled from Python to ST codes using GPT-4, with filtering applied to exclude examples that were either too complex algorithmically or too specific to Python. This conversion process resulted in over 2000 ST code samples, which were subsequently used for SFT. During the experiments, Parameter-Efficient Fine-Tuning (PEFT) [28] was employed as the fine-tuning approach, with Low-Rank Adaptation (LoRA) [15] used to fine-tune the language model via DPO. This configuration was chosen to optimize performance by minimizing the number of trainable parameters, significantly reducing computational resource consumption. The LoRA modules were integrated into the transformer layers of the model, enabling efficient adaptation while preserving the underlying pre-trained knowledge and avoiding significant deviation from the baseline model.
Due to the limited availability of compilers compliant with the IEC 61131-3 standard [3] for the ST language, the semantic expert was selected based on the capabilities of the recently developed open-source compiler, Rusty [1]. Rusty offers the essential features required for syntactical analysis and facilitated the integration of this expert into the overall framework. For the semantic expert, a state-of-the-art model from OpenAI was chosen. Dedicated test runs were conducted to quantify the added value of GPT-3.5 Turbo as a code generator and GPT-4 as a semantic expert. This setup allowed for cross-validation between different GPT versions, ensuring that one model was not evaluating its own performance. Both models were provided with the same expert prompt (cf. Fig. 2), incorporating additional contextual elements to enhance their respective pre-existing knowledge.
@@
[…]
@@ IV-C Results and Discussion
The graphs in Fig. 3 break down the improvements across training iterations. Iteration 0 represents the ”out-of-the-box” performance of the model, while Iteration 1 reflects the results after applying SFT. From Iteration 2 to 11, we observe the performance following the application of DPO. As depicted, there is a notable upward trend in the performance metrics, with the most significant improvements occurring after Iteration 2. After that, the metrics plateau, indicating stabilization in model performance.
@@
As illustrated in Fig. 3A, the results demonstrate a significant enhancement in the compilation success rate compared to the baseline model, denoted as $\psi$ (e.g., the SFT fine-tuned Phi-3). The proposed method achieves a compilation success rate of 70%, a substantial increase from the baseline’s 5%. This also surpasses the performance of GPT-3.5, which yielded a 40% success rate. Although feedback in DPO is provided in binary form rather than through multiple reward values based on compilation status, the DPO loss function effectively compensates for this by utilizing its internal reward mechanism.
@@
As shown in Fig. 3B, the model’s semantic correctness rate peaked at around 45%. This improvement was facilitated by the integration of a dedicated LLM-based expert $\mathcal{E}_{\phi}$, which enabled the implementation of a ranking policy through its textual comparison capabilities. The expert’s ability to evaluate each code sample $\hat{C}_{i}$ contributed significantly to the refinement of the model and the overall training process. This suggests that DPO is not only effective in natural language tasks but is also highly efficient in programming code generation that requires strict adherence to syntactic and semantic rules.
@@
The joint compilation and semantic correctness rate depicted in Fig. 3C of approximately 39% was achieved after several iterations. This integrated metric highlights the reliability and applicability of the generated ST code in industrial automation, where both accuracy and functionality are essential. The iterative process of training, evaluation, and feedback was critical to achieving these results.
@@
Overall, substantial improvements have been demonstrated when comparing this approach to baseline models. The improvements observed are primarily due to the enhanced quality of training data via robust validation and overall domain-specific adaptation.
@@
This suggests that further refining feedback loops and data collection processes could lead to even greater performance gains. One potential area for further development involves the use of stronger baseline models. While Phi-3 14B performs well in many natural language processing and reasoning tasks, it may encounter challenges with more advanced logic, complex algorithms, or generalization to intricate intents. Future exploration could address these challenges, improving its ability to handle complex tasks. LoRA efficiently fine-tunes models using fewer parameters, but this reduced parameter count may limit the model’s ability to capture the full depth of certain programming languages, especially those it hasn’t been deeply trained on. Future efforts could combine LoRA with other fine-tuning techniques to better capture the complex syntactic and contextual details needed for high-quality code generation. When evaluating code quality, the model relies on patterns learned from training data, which may not always cover edge cases or specific application needs. This emphasizes the importance of incorporating software testing methods, such as unit and functional testing, to complement the model’s evaluations.
@@
Finally, the present work considers only PLC programming in the ST language; however, it will be interesting to explore Structured Control Language (SCL) due to its unique syntactic differences that could benefit from further refinement.
@@ V Conclusion
The methodology presented in this paper employs DPO to improve code generation for PLC programming using ST. By integrating compiler feedback and expert semantic evaluations, the approach iteratively refines the quality of the training data, leading to more accurate and reliable code generation. The use of synthetic data and a preference-based fine-tuning process specifically addresses data scarcity issues, allowing the model to perform better without modifying its architecture. The significant improvements observed over baseline models are largely attributed to the enhanced training data, highlighting the potential of this pipeline for efficiently automating PLC programming in industrial environments.
The experimental results validate the effectiveness of this DPO-driven pipeline in generating precise and reliable PLC Structured Text code. The substantial improvements in compilation success, semantic correctness, and joint performance metrics highlight the power of advanced language models, fine-tuned with domain-specific data and reinforced through iterative feedback processes. These advancements indicate a promising direction for the automation and optimization of PLC programming, paving the way for more scalable, reliable, and efficient industrial automation solutions.

### v3

@@ IV-A Performance Metrics
To quantify the performance of the method, we define evaluation metrics appropriate for the ST experts which are abbreviated from [31][19].
@@
    Compilation rate, $P_{c}$ – The empirical probability that a code segment $C_{i}$ successfully compiles. This metric reflects the likelihood that the code passes syntax and type-checking stages without errors.
@@
    Semantic correctness rate, $P_{s}$ – The empirical probability that $C_{i}$ is semantically accurate. The primary criterion is the quality of the implementation and its alignment with the intent.
@@
    Joint semantic correctness and compilation rate, $P_{j}$ – The empirical joint probability that a generated code sample is both syntactically valid and semantically correct. This metric reflects the overall success in passing both compilation and semantic validation processes.
These customized metrics provide evaluation for the quantitative aspects of the method.
SYSTEM: “You are an expert in PLC Structured Text (ST) programming as per IEC 61131-3 standards. Your task is to evaluate the given code based on its logical, semantic, and syntactic correctness. You can tolerate minor mistakes or errors. IMPORTANT: Use the following format for your responses: “[0] [1] [1]”, where 0 indicates a negative evaluation and 1 indicates a positive evaluation. IMPORTANT: Ensure that the number of brackets matches the number of given implementations. IMPORTANT: Provide the evaluations strictly in the specified format without any additional text, code, or formatting. INSTRUCTION: Given implementations are separated by ’====================’.” USER: “Great! Here is another problem and its corresponding implementations to evaluate. Problem: $X_{i,n}$ Implementations: subset of $Y_{i}$ that compiles or $Y_{i}$ if no samples compile”
Figure 2: The expert prompt consists contains the specific intent $X_{i,n}$, along with generated code samples from $Y_{i}$
and an ST expert instruction.
@@ IV-B Experimental Setup
Parameter Description
[…]
Selection Strategy $\pi_{i}\sim\mathcal{U}\left(\frac{1}{M}\right)\forall i\in\{1,\dots,I\}$
Supervised Dataset ($\mathcal{D}_{SFT}$) 300 OSCAT, 1000 converted APPS
Instruction Set ($\mathcal{X}$) 100 OSCAT, 900 converted APPS intents
Number of code responses T = 15
Iterations I = 9
Evaluation Dataset 100 OSCAT, 100 converted APPS intents
Semantic ST Expert gpt-3.5-turbo-1106 + prompt (Fig. 2)
Syntactic ST Expert Rusty compiler [33]
Evaluators Rusty compiler + gpt-4-1106-preview
Table I: Experimental Parameters for the Study
The baseline model $\theta$ is based on the Phi-3 LLM architecture [34], specifically the version with 14 billion parameters (Phi-3 14B). To maintain its optimal performance, no quantization techniques were applied. The model was reparameterized using SFT with a synthetic dataset generated in advance. This dataset was derived from the APPS dataset [10], which contains prompt-code pairs. The implementations were transpiled from Python to ST codes using GPT-4, with filtering applied to exclude examples that were either too complex algorithmically or too specific to Python. This conversion process resulted in ca. 2000 ST code samples, which were subsequently used for SFT. During the experiments, Parameter-Efficient Fine-Tuning (PEFT) [35] was employed as the fine-tuning approach, with Low-Rank Adaptation (LoRA) [36] used to fine-tune the language model via DPO. This configuration was chosen to optimize performance by minimizing the number of trainable parameters, significantly reducing computational resource consumption. The LoRA modules were integrated into the transformer layers of the model, enabling efficient adaptation while preserving the underlying pre-trained knowledge and avoiding significant deviation from the baseline model.
@@
IV-C IEC 61131-3 ST Experts
There is a limited availability of compilers compliant with the IEC 61131-3 standard [37] for the ST language. The syntactic expert is selected based on the capabilities of the open-source compiler, Rusty [33]. Rusty offers the essential features required for syntactical analysis and facilitated the integration of this expert into the overall framework. This expert returns a binary feedback on the compilation success. For the semantic expert ChatGPT-3.5 is used and queried for a binary evaluation of given ST code. For the evaluation as shown in 3 the semantic expert is GPT-4. This setup aims to allow for cross-validation between different GPT versions, ensuring that one model was not evaluating its own performance. Both models were provided with the same expert prompt Fig. 2, incorporating an evaluation prompt and corresponding original intent for the generated code to be evaluated.
2222444466668888000.10.10.10.10.30.30.30.30.50.50.50.50.70.70.70.7Iteration $i$$P_{c}$
2222444466668888000.10.10.10.10.30.30.30.30.50.50.50.50.70.70.70.7Iteration $i$$P_{s}$
2222444466668888000.10.10.10.10.30.30.30.30.50.50.50.50.70.70.70.7Iteration $i$$P_{j}$
@@ IV-D Results and Discussion
The graphs in Fig. 3 break down the improvements across number of $I=9$ training iterations. Also, the performance of the base model $\theta$, as well as the STF variant $\theta_{\text{sft}}$ is shown. As a closed-source competitor ChatGPT-3.5 is shown. From Iteration 1 to 9, we observe the performance following the application of DPO. As depicted, there is a notable upward trend in the performance metrics, with the most significant improvements occurring after Iteration 2. After that, the metrics plateau, indicating no further expected improvements.
@@
As illustrated in Fig. 3A, the results demonstrate a significant enhancement in the compilation success rate compared to the baseline model. The proposed method achieves a compilation success rate of around $70$%, a substantial increase from the baseline’s 7%. This also surpasses the performance of GPT-3.5, which yielded a 40% compile rate. Although feedback in DPO is provided in binary form rather than through multiple reward values based on compilation status, it seems like the DPO loss function effectively compensates for this by utilizing its internal reward mechanism.
@@
As shown in Fig. 3B, the model’s semantic correctness rate peaked at around 45%. This improvement was facilitated by the integration of semantic expert represented by ChatGPT-3.5, which enabled the implementation of a ranking policy through its textual comparison capabilities. The expert’s ability to evaluate each code sample $\hat{C}_{i}$ contributed to the refinement of the model and increasement in both syntactic and semantic correctness. This suggests that DPO is not only effective in natural language tasks but is also highly efficient in programming code generation that requires strict adherence to syntactic and semantic rules.
@@
The joint compilation and semantic correctness rate depicted in Fig. 3C of approximately 39% was achieved after several iterations. This metric highlights the reliability and applicability of the generated ST code in industrial automation, where both accuracy and functionality are essential. The iterative process of training, evaluation, and feedback was critical to achieving these results.
@@
Overall, substantial improvements have been demonstrated when comparing this approach to baseline models. The improvements are shown to be relatively constantly growing over iterations.
@@
This suggests that further refining feedback loops and data collection processes could lead to better performance gains. One potential area for further development involves the use of stronger baseline models. While Phi-3 14B performs well in many natural language processing and reasoning tasks, it may encounter challenges with more advanced logic, complex algorithms, or generalization to intricate intents. Future exploration could address these challenges, improving its ability to handle complex tasks. LoRA efficiently fine-tunes models using fewer parameters, but this reduced parameter count may limit the model’s ability to capture the full depth of certain programming languages, especially those it hasn’t been deeply trained on. Future efforts could combine LoRA with other fine-tuning techniques to better capture the complex syntactic and contextual details needed for high-quality code generation. When evaluating code quality, the model relies on patterns learned from training data, which may not always cover edge cases or specific application needs. This emphasizes the importance of incorporating software testing methods, such as unit and functional testing, to complement the model’s evaluations.
@@
Finally, the present work considers only PLC programming in the ST language; however, it will be interesting to explore Structured Text (ST) due to its unique syntactic differences that could benefit from further refinement.
@@ V Conclusion
The methodology presented in this paper employs an iterative approach to fine-tune a LLM for generation of Structured Text by using a supervised-fine-tuning process on a small dataset and iterative DPO based on a compiler and LLM expert. The integration of the compiler feedback and the LLM expert in form of a semantic correctness feedback leads to better results than the SFT model. The use of synthetic data and a preference-based fine-tuning process specifically addresses the data scarcity associated with domain-specific programming languages, making it possible to train LLMs without extensive data collection or costly human labeling steps. This approach enables the development of industry-specific copilots that can revolutionize how PLCs are programmed, paving the way for total automation of industry.